### v2

@@ 3.3 The Impact of Peers on Cadets’ Allegiance
Additionally, consistent with Figure 2, we find that peer influence is not significant for cadets from Free States (Columns (4)-(6)). Based on these findings, our subsequent analysis focuses on cadets from Slave States.
Peer Effects by Type of State
As we see in the analysis above, there is no peer influence on cadets in the Free States. We can further divide the Slave States into groups by the relative slave populations. It turns out that the peer influence operates entirely on cadets from the middle range of Slave States that have slave populations below one third of the population.
@@
In Table 2, we break states into three groups: those without slaves, those with slaves up to one third of the population, and those with slave populations above one third. We now see that the peer effect comes entirely from this middle group of states, and in fact the estimated coefficient and significance are larger than that from the previous table.
These findings suggest that in High-Slavery States, economic imperatives and entrenched social norms overrode peer influence, muting its effect.
Table 2: Peer composition and allegiance choice: different groups of states
@@
Share Free-State peers (sd) 0.009 0.062∗∗ -0.005
@@
Share Border-State peers (sd) -0.027 0.022 -0.021
@@
Dependent var.mean 0.115 0.476 0.919
Observations 122 267 540
R-squared 0.168 0.190 0.032
@@
Note. The sample for this analysis consists of all war-participating cadets. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
3.4 Career Outcomes
We know the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their life results, such as military rank and survival probability. We do not presume that this had any influence on their allegiances, but does trace some consequences of their decisions.
As indicated in Columns (1)-(2) of Table 3, joining the Union correlates with a lower military rank in 1865 and a reduced probability of attaining a rank of General-level positions. This aligns with the fact that the Confederate army, was both smaller and had more generals per enlisted man, and starting from scratch had no existing officers, enabling quicker promotions and a higher likelihood of becoming a general for people with officer training. On the other hand, joining the Union is associated with a lower risk of dying in the Civil War, reflecting the higher fatality rate on the Confederacy side, as reported in Column (3).
Table 3: Subsequent outcomes in 1865
OLS IV
Rank General Died Rank General Died
@@ 3.4 Career Outcomes
A. Slave-State cadets (>1%)
Joining the Union -1.556∗∗∗ -0.346∗∗∗ -0.168∗∗∗ -1.952∗∗∗ -0.369∗∗∗ -0.137
@@ A. Slave-State cadets (>1%)
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 6.974 0.435 0.207 6.974 0.435 0.207
Observations 382 382 389 382 382 389
F-statistic 9.068 9.068 9.375
R-squared 0.314 0.187 0.095
B. Slave-State cadets (1%-33%)
Joining the Union -1.510∗∗∗ -0.318∗∗∗ -0.171∗∗∗ -1.870∗∗∗ -0.373∗∗ -0.161
@@ B. Slave-State cadets (1%-33%)
Controls Y Y Y Y Y Y
State FEs Y Y Y Y Y Y
Dependent var. mean 6.784 0.402 0.184 6.784 0.402 0.184
Observations 264 264 267 264 264 267
F-statistic 8.133 8.133 8.365
R-squared 0.310 0.187 0.081
Note. The sample for Panel A consists of the war-participating cadets from Slave States (i.e., with a slave population higher than 1%), and the sample for Panel B excludes those from the High-Slave States who were not significantly influenced by peers. Columns (1)-(3) use OLS estimation, while Columns (4)–(6) present IV estimates using birthplace indicators and the share of peers from Free States as instruments for Union affiliation. Columns (1) and (4) feature the military rank as the dependent variable, classified into 11 levels, with General assigned 11 and Third Lieutenant assigned 1. General and Died are binary variables. The rank information on seven cadets is unknown. The standard errors reported in the tables are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
3.5 Peer Influence and Interaction
Our baseline results indicate that peers shaped individual decisions at a pivotal moment. To shed light on the mechanisms behind this influence, we focus on two complementary dimensions of interaction. First, we test whether peer effects are stronger within a cadet’s own cohort compared to across adjacent classes. Second, we assess whether the strength of influence depends on shared military experience: in particular joint participation in the Mexican-American War.
Influence by Cohort
We analyze peers by grouping them into cohorts and display the coefficients for peers within the same cohort, as well as for cohorts from t-1 to t-2 and t+1 to t+2, in Supplemental Appendix Figure S3. The coefficients in the figure are obtained from a single regression that includes all controls from our baseline analysis. The result reveals that peer influence is most important within the same cohort. In contrast, although there is some degree of influence from earlier cohorts (older cadets), future cohorts do not show comparable levels of influence (younger cadets).
Relevance of Co-fighting in the Mexican-American War
A significant share—34.6%—of cadets in our dataset participated in the Mexican-American War (1846–1848), offering a valuable context to examine how shared experiences shape peer influence. For each cadet, we compute the proportion of peers who also fought in the war, distinguishing between those from slave and Free States.
If shared wartime experience intensifies peer effects, we would expect our core finding on peer influence to be stronger when a cadet had more co-fighting peers from Free States, and weaker when more such peers were from Slave States. This is precisely what we observe (see Supplemental Appendix Table S9), focusing cohorts between 1820 and 1845. Although the limited sample size prevents precise estimation of this interaction effect, the magnitude is economically meaningful, suggesting that shared combat experience enhances the intensity of peer influence.
@@ 3.6 Additional Analyses
We conduct five supplementary analyses to account for (1) whether the effect of peers on cadets could vary with shifting political environments, (2) potential differences between peers who graduated and those who dropped out, (3) whether non-participants in the war affect the results, (4) whether cash-crop employment shapes peer influence, and (5) whether defining peers by appointment state rather than home state matters. We describe the results from these analyses below.
@@
First, we divide the sample into periods before and after 1850. Although the issue of slavery was highly contentious throughout the period we study, it was particularly intense in the 1850s and nearly led to the collapse of the United States. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S10, peer influence (though estimated imprecisely due to limited sample sizes) persisted over time, with similar effects on either side of this divide.
@@
As discussed above, peer composition is not systematically correlated with one’s choice of participating in the war or not. Thus, we have focused on joining the Union versus the Confederacy for simplicity. If we further employ a multinomial logit model to consider whether a cadet joined the Union, joined the Confederacy or did not participate in the war, we again find that peer composition is not predictive for whether a cadet participated in the war, while our main finding of peer influence on joining the Union vs Confederacy holds. These results are presented in Supplemental Appendix Table S13. According to these estimates on the log of the odds ratios, a one-standard deviation increase in Free-State peers increased probability of joining the Union by approximately 4.4 percentage points, higher than our baseline estimates.
@@
Unlike the results based on home-state slave share, we find no clear interaction between peer influence and cash-crop employment history (Supplemental Appendix Table S15). This may be due in part to the relatively small proportion—only 7%—of cadets from Slave States having such employment experience.
@@ 4 Discussion
Our study also provides a road map for identifying such peer effects in other high-stakes and high-polarization settings. This sort of peer influence is challenging to quantify and identify causally through anecdotal evidence alone. It is also not enough to identify correlation in people’s behaviors, given that friendships are endogenous and people are subject to common (potentially unobserved) factors (Angrist, 2014). West Point’s environment—characterized by structured, close-knit interactions among cadets from diverse geographic backgrounds—provided us an opportunity to analyze the influence of quasi-random peer exposure, and its interplay with background allegiances. By using variation in a cadet’s peers’ state-of-origin for identification, we have something that is exogenous to other key things influencing a given cadet’s decision. Similar techniques can be used in other high-stakes contexts.
That said, our ability to definitively trace the mechanisms behind peer influence is limited by the nature of our historical data. Our specification that examines cadets’ choices based on the composition of their cohort by the states of origin of their peers enables us to causally identify peer influence. Nonetheless, it does not provide insight into why peers influence cadets’ decisions. One curious aspect is that the peer influence only affected cadets from states with slaves. This could be due to the fact that West Point was an institution in which cadets were being trained for the U.S. Army, and so only cadets from Slave States faced a conflicting allegiance. Our evidence on the strength of within-cohort influence and how joint participation in the Mexican-American War amplifies peer influence suggests that shared-experiences mattered.

### v3

@@ 3.3 The Impact of Peers on Cadets’ Allegiance
In contrast to the strong peer influence observed among cadets from Slave States, we find no comparable peer effect among cadets from Free States, consistent with Figure 2. These results are reported in Columns (5)–(8).
Next, we interpret this asymmetric peer effect in a simple framework.
Interpreting the Asymmetric Peer Effects
The empirical asymmetry in peer influence—namely that cadets from Slave States respond to peer composition while cadets from Free States do not—can be understood through a simple choice framework.
Normalize the value of joining the Confederacy to zero. A cadet joins the Union if the net utility from doing so is positive. This utility consists of three components. First, there is a default institutional allegiance associated with West Point and service in the U.S. Army. Let this term be $V>0$. Second, cadets from Slave or Border States may have material or ideological interests in the Confederacy—denoted $S_{i}\geq 0$, with $S_{i}=0$ for cadets from Free States. Third, peer interactions depend on the fraction of a cadet’s cohort originating from Free States, $F_{i}$. We capture peer influence via $\beta(F_{i}-1/2)$, which increases (decreases) the utility of joining the Union when Free-State peers are in the majority (minority).
Total utility of joining the Union for cadet $i$ is
$U_{i}=V-S_{i}+\beta(F_{i}-1/2)+\varepsilon_{i},$
where $\varepsilon_{i}$ is an idiosyncratic shock with distribution $G(\cdot)$.
Cadets thus join the Union as long as
$\varepsilon_{i}>S_{i}-V-\beta(F_{i}-1/2).$
Thus, the probability of joining the Union is for cadet $i$ is:
$P_{i}=1-G(S_{i}-V-\beta(F_{i}-1/2)).$ (2)
For Free-State cadets $S_{i}$ and $V$ is likely to be high, so this probability is close to 1. For cadets heavily exposed to slavery (either from a region heavily dependent on slavery, or personally owning slaves), $S_{i}$ is large and this probability is close to 0. It is for intermediate values of $S_{i}-V$ that this varies substantially as a function of $\beta F_{i}$.
For example, if $G$ is a uniform distribution on $[0,1]$, then (2) becomes
$P_{i}={\rm min}\left[1,\left(1-\beta/2+V-S_{i}+\beta F_{i}\right)^{+}\right],$
which is a constant 1 for low $S_{i}$ and $V>1-\beta/2$ (Free-State cadets), consistent with what we have found above. It is a constant 0 for high $S_{i}$, and has a slope of $\beta$ in $F_{i}$ for intermediate values of $S_{i}$. Below, we verify this interaction of the peer effect with the magnitude of the slave interest of cadets ($S_{i}$).
More generally, differentiating (2) yields
$\frac{\partial P_{i}}{\partial F_{i}}=\beta\,g\!\left(S_{i}-V-\beta(F_{i}-1/2)\right),$
where $g(\cdot)$ is the density of $G(\cdot)$.
For Free-State cadets ($S_{i}=0$) and sufficiently large $V$, this is close to zero because choices are far from the margin where the density is small, consistent with what we found above. Moreover, among Slave-State cadets, a high $S_{i}$ mitigates peer effects. When $S_{i}$ is very large, $\frac{\partial P_{i}}{\partial F_{i}}$ is also far from the margin and close to zero. With this in hand, we examine how peer influence varies with various proxies for $S_{i}$.
Variations in Peer Effects with Political-Economic Backgrounds
We use several proxies to measure cadets’ material and ideological interests in supporting the Confederacy. Our main focus is on exposure to and personal investment in slavery: the slave population share in a cadet’s home state, the slave population share in his home county, and whether he owned slaves in 1860. In Panel A of Table 2, we interact our peer measure with the state-level slave population share and with cadet slave ownership. The peer effect is smaller where slavery was more prevalent, and likewise among cadets who owned slaves. Deeper entanglement in the slave economy reduced responsiveness to peers. Correspondingly, once we control for this the magnitude of the peer effect among the remaining slave-state cadets becomes even larger. For example, among slave-state cadets who did not personally own slaves, we see that the coefficient is now .076, more than a third increase over the coefficient not controlling for this.
@@
We further divide the Slave States into groups by the relative slave populations. It turns out that the peer influence operates entirely on cadets from the middle range of Slave States that have slave populations below one third of the population. In Panel B of Table 2, we break states into three groups: those without slaves, those with slaves up to one third of the population, and those with slave populations above one third. We now see that the peer effect comes entirely from this middle group of states, and in fact the estimated coefficient and significance are larger than that from the previous table. These findings suggest that in High-Slavery States, economic imperatives and entrenched social norms overrode peer influence, muting its effect. This is consistent with a very high $S_{i}$ in equation (2).
Table 2: Peer composition and allegiance choice: different groups
@@
State level County level Individual level
(1) (2) (3)
A. Interaction with Slave Variables
Share Free-State Peers (sd) 0.053∗∗∗ 0.053∗∗ 0.076∗∗∗
(0.019) (0.021) (0.023)
Share Free-State Peers (sd) $\times$ Slave Pop. Share (State) -0.030
Share Free-State Peers (sd) $\times$ Slave Pop. Share (County) -0.036
Share Free-State Peers (sd) $\times$ Cadet Slave Ownership -0.083∗∗
Controls Y Y Y
State FEs Y Y Y
Dependent var.mean 0.361 0.358 0.361
Observations 388 352 388
R-squared 0.291 0.341 0.357
B. Different Groups of States Join the Union: War Participants
@@
Share Free-State Peers (sd) 0.010 0.072∗∗ -0.013
@@
Share Border-State Peers (sd) -0.044 0.009 -0.018
@@
Dependent var.mean 0.107 0.477 0.919
Observations 122 266 540
R-squared 0.149 0.205 0.033
@@
Note. The sample for this analysis consists of all war-participating cadets. The control variables consist of Age in 1860, Class Rank, and Cohort. For Panel A, Columns (1)–(3) include controls for state-level Slave Population Share (sd), county-level Slave Population Share (sd), and Cadet Slave Ownership, respectively. The standard errors reported in the parentheses are derived from bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
Religious and political preferences could be relevant and are also correlated with slave population share (Fogel, 1994). To explore this potential influence, we incorporate county-level measures of religion and voting behavior, drawing on census data on church seats by denomination and historical election returns. We focus this analysis on Slave States, as those are where the peer effects operated.
For religion, we use the number of (Southern) Baptist churches per 1,000 capita as a proxy for a more pro-slavery religious environment, and the number of Friends (Quaker) churches per 1,000 capita as a proxy for an antislavery religious environment. Prior researchers have noted that among the Protestant denominations, (Southern) Baptists were generally more supportive of slave ownership. This followed the split of the Baptist Church in 1845, which produced the Southern Baptist Convention after national mission boards refused to appoint slaveholders as missionaries. This has been interpreted by historians as an indication of (Southern) Baptist support for the legitimacy of slave holding prior to the Civil War (Mathews, 1977; Noll, 2006). Although the split was less clear in the border states—where Northern and Southern affiliations coexisted—Baptists in these regions still tended to oppose abolitionism (Harlow, 2014). Other Protestant churches (e.g., Methodists and Presbyterians) had more diverse views regarding slavery (Wigger, 2001; Snay, 1993).
For voting, we use the vote share for the Southern Democratic candidate (Breckinridge) as a measure of local political support for slavery, and the vote share for the Republican candidate (Lincoln) in Slave States as a measure of local political opposition to slavery.
As shown in Appendix Table S9, including these variables and their interactions with peer exposure does not alter our main findings concerning peer influence among cadets from Slave States nor the interaction of the peer effect with slave population share. Among the newly added controls, the antislavery vote share (Lincoln) is positively associated with the probability of joining the Union, consistent with expectations.
These results also suggest that Slave Population Share and Cadet Slave Ownership are more powerful proxies for the most relevant economic, social, and political considerations than these other proxies.
3.4 Shared Experience and Peer Influence
Our baseline results indicate that peers shaped individual decisions at a pivotal moment. To shed light on the mechanisms behind this influence, we focus on three complementary dimensions of interaction. First, we test whether peer effects are stronger within a cadet’s own cohort compared to across adjacent classes. Second, we distinguish between cadets who remained in continuous military service and those who temporarily exited the Army before the Civil War, allowing us to gauge whether sustained professional identity amplified peer influence. Third, we assess whether the strength of influence depends on shared military experience: in particular joint participation in the Mexican-American War. Together, they reveal the importance of shared experience behind peer influence.
Influence by Cohort
To examine how cohort exposure matters, we analyze peers by grouping them into cohorts and display the coefficients for peers within the same cohort, as well as for cohorts from t-1 to t-3 and t+1 to t+3, in Supplemental Appendix Figure S3. The coefficients in the figure are obtained from a single regression that includes all controls from our baseline analysis. The result reveals that peer influence is most important within the same cohort. In contrast, although there is some degree of influence from earlier cohorts (older cadets), future cohorts do not show comparable levels of influence (younger cadets).
Non-continuous vs. Continuous Military Service
Distinguishing between cadets who had returned to civilian life between graduation and 1860 and those who remained professional soldiers during that time can provide further understanding of the operation of peer effects.
To do this, we manually reviewed military service records and coded whether and when individuals exited the regular army prior to the Civil War. A key complication is that all cadets who joined the Confederacy were required to resign from the U.S. Army, so treating any exit as a return to civilian life would mechanically misclassify the Confederates. To avoid this, we conservatively classify a cadet as having returned to civilian life only if he exited the regular army and remained out of military service for more than one year before the start of the war. Using this definition, we split the sample into cadets with non-continuous versus continuous military service.
As shown in Appendix Table S10, peer exposure is positive in both groups but is substantially larger and statistically significant only among cadets with continuous military service, suggesting an important role for shared professional military experience.
Relevance of Co-fighting in the Mexican-American War
A significant share—26.1%—of cadets in our dataset participated in the Mexican-American War (1846–1848), offering a valuable context to examine how shared experiences shape peer influence. For each cadet, we compute the proportion of peers who also fought in the war, distinguishing between those from slave and Free States.
If shared wartime experience intensifies peer effects, we would expect our core finding on peer influence to be stronger when a cadet had more co-fighting peers from Free States, and weaker when more such peers were from Slave States. This is precisely what we observe (see Table 3), focusing on cohorts between 1820 and 1845. Although the limited sample size prevents precise estimation of this interaction effect, the magnitude is economically meaningful, suggesting that shared combat experience enhances the intensity of peer influence.
Table 3: Peer composition and allegiance choice: Mexican-American War
Dependent var Join the Union: War Participants
Slave States
Slave-State Cadets (>1%) Slave-State Cadets (1%-33%)
@@ 3.4 Shared Experience and Peer Influence
Sh. Free-State Peers join Mex.-Am. War $\times$ Sh. Free-State Peers 0.160 0.117 0.109 0.220∗∗ 0.201 0.212
@@
Sh. Slave-State Peers join Mex.-Am. War $\times$ Sh. Free-State Peers -0.114 -0.084 -0.068 -0.172 -0.155 -0.156
@@
Share Free-State Peers (sd) 0.059 0.044 0.041 0.050 0.034 0.025
(0.036) (0.033) (0.035) (0.045) (0.046) (0.048)
Share Free-State Peers join Mex.-Am. War (sd) -0.109 -0.066 -0.058 -0.156 -0.117 -0.148
(0.086) (0.079) (0.095) (0.104) (0.106) (0.120)
Share Slave-State Peers join Mex.-Am. War (sd) 0.091 0.054 0.058 0.140 0.113 0.147
(0.084) (0.079) (0.092) (0.099) (0.099) (0.109)
Controls N Y Y N Y Y
State FEs N N Y N N Y
Dependent var. mean 0.402 0.402 0.402 0.515 0.515 0.515
Observations 184 184 184 136 136 136
R-squared 0.035 0.264 0.323 0.047 0.163 0.233
Note. This table presents the impact of the fraction of peers from Free States in a cadet’s cohort on that cadet’s decision to join the Union. Share Free-State peers join Mexican-American War: The proportion of Free-State peers participating in Mexican-American war with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. Share Slave-State peers joining the Mexican-American war: The proportion of Slave-State peers participating in Mexican-American War with themselves. If the cadet did not participate in the Mexican-American War, the proportion is 0. The control variables consist of Age in 1860, Class Rank, state level Slave Population Share (sd) and Cohort. The standard errors presented in the parentheses are obtained through bootstrapping with 400 resampling iterations. ***$p<0.01$, **$p<0.05$, *$p<0.1$.
@@ 3.5 Additional Analyses
We conduct six supplementary analyses to account for (1) whether the effect of peers on cadets could vary with shifting political environments, (2) potential differences between peers who graduated and those who dropped out, (3) whether non-participants in the war affect the results, (4) whether cash-crop employment shapes peer influence, and (5) whether defining peers by appointment state rather than home state matters. Finally, in our exercise (6), we examine post-war career outcomes to trace some consequences of cadets’ choices.
@@
First, we divide the sample into periods before and after 1850. Although the issue of slavery was highly contentious throughout the period we study, it was particularly intense in the 1850s and nearly led to the collapse of the United States. In an effort to avert civil war, Congress passed the Compromise of 1850, which included the Fugitive Slave Law that obligated law enforcement nationwide to assist in capturing alleged runaway slaves, and many in the north resisted its enforcement (McPherson, 2003). Against this backdrop, we use 1850 as a dividing point to assess whether peer influence experienced notable changes before and after that year. As illustrated in Supplemental Appendix Table S11, peer influence persisted over time, with similar effects on either side of this divide.
@@
As discussed above, peer composition is not systematically correlated with one’s choice of participating in the war or not. Thus, we have focused on joining the Union versus the Confederacy for simplicity. If we further employ a multinomial logit model to consider whether a cadet joined the Union, joined the Confederacy or did not participate in the war, we again find that peer composition is not predictive for whether a cadet participated in the war, while our main finding of peer influence on joining the Union vs Confederacy holds. These results are presented in Supplemental Appendix Table S14. According to these estimates on the log of the odds ratios, a one-standard deviation increase in Free-State peers increased probability of joining the Union by approximately 7 percentage points, higher than our baseline estimates.
@@
Unlike the results based on slavery proxies, we find no clear interaction between peer influence and cash-crop employment history (Supplemental Appendix Table S15, Columns (4) and (8)). This may be due in part to the relatively small proportion—only 4%—of cadets from Slave States having such employment experience.
@@
Career Outcomes
We know the post-war outcomes of many who joined an army, allowing us to investigate how their decisions during this critical historical juncture influenced their ex post life results, such as military rank and survival probability. We do not presume that this had any influence on their allegiances, but does trace some consequences of their decisions.
As indicated in Columns (1)-(2) of Supplemental Appendix Table S17, joining the Union correlates with a lower military rank in 1865 and a reduced probability of attaining a rank of General-level positions. This aligns with the fact that the Confederate army, was both smaller and had more generals per enlisted man, and starting from scratch had no existing officers, enabling quicker promotions and a higher likelihood of becoming a general for people with officer training. On the other hand, joining the Union is associated with a lower risk of dying in the Civil War, reflecting the higher fatality rate on the Confederacy side, as reported in Column (3).
@@ 4 Discussion
Our study also provides a possibility for identifying such peer effects in other high-stakes and high-polarization settings. This sort of peer influence is challenging to quantify and identify causally through anecdotal evidence alone. It is also not enough to identify correlation in people’s behaviors, given that friendships are endogenous and people are subject to common (potentially unobserved) factors (Angrist, 2014). West Point’s environment—characterized by structured, close-knit interactions among cadets from diverse geographic backgrounds—provided us an opportunity to analyze the influence of quasi-random peer exposure, and its interplay with background allegiances. By using variation in a cadet’s peers’ state-of-origin for identification, we have something that is exogenous to other key things influencing a given cadet’s decision. Similar techniques can be used in other high-stakes contexts.
That said, our ability to definitively trace the mechanisms behind peer influence is limited by the nature of our historical data. Our specification that examines cadets’ choices based on the composition of their cohort by the states of origin of their peers enables us to causally identify peer influence. Nonetheless, it does not provide insight into why peers influence cadets’ decisions.
We provide a simple model that can indicate why the peer influence only affected cadets from states with slaves, and primarily among those cadets who did not own slaves nor lived in regions heavily dependent upon slaves. Our evidence on the strength of within-cohort influence, the relevance of continuous military experience, and how joint participation in the Mexican-American War amplifies peer influence suggests that shared-experiences mattered.